OpenStudy (anonymous):

What is the solution to equation, square root of (2x+8) -6=4

OpenStudy (anonymous):

Helppppp

OpenStudy (anonymous):

Does it look like this? \[\large \sqrt{2x+8}-6=4\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Well the first step is to get the 6 to the other side. Can you do that?

OpenStudy (anonymous):

yes you add it to the 4 to get 10. but I have no idea what to do after that

OpenStudy (anonymous):

So now it will look like this \[\large \sqrt{2x+8}=10\] So we're going to get x out of the square root by squaring both sides. So \[\large \sqrt{2x+8}^{2}=10^{2}\]

OpenStudy (anonymous):

so is it going to be, (2x+8)(2x+8)=20?

OpenStudy (anonymous):

No. \[\large \sqrt{x}^{2} = x\] So \[\large \sqrt{2x+8}^{2} = 2x+8\] And \[\large 10^{2}=10 \times 10\]

OpenStudy (anonymous):

So 2x+8=100. and then you get ride of the 8 by subtracting on the other side and the divide by 2. would the answe be x=46?

OpenStudy (anonymous):

Yes, that would be it. Good job :)

OpenStudy (anonymous):

Thank you!! (:

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]